While running, a 78 kg student generates thermal energy at a rate of 1100 W . To maintain a constant body temperature of 40 ∘C, this energy must be removed by perspiration or other mechanisms.

If these mechanisms failed and the heat could not flow out of the student's body, for what amount of time could a student run before irreversible body damage occurred? (Protein structures in the body are damaged irreversibly if the body temperature rises to 44 ∘C or above. The specific heat capacity of a typical human body is 3480 J/(kg⋅K), slightly less than that of water. The difference is due to the presence of protein, fat, and minerals, which have lower specific heats.)
Answer #1

given :

rate of thermal energy generation in body, P = 1100 W (or Joule/sec)

mass of body, m = 78 kg

initial body temperature, Ti = 40 C

final body temperature, Tf = 44 C

specific heat capacity of body, C = 3480 J/kg-K

hence,

increase of internal energy of body, U = m*C* = m*C*(Tf - Ti)

U = 78*3480*(44 - 40)

U = 1085760 Joule

hence, time taken by body to increase temperature from 40 C to 44 C